# Exploration and Incentivizing Participation in Clinical Trials

### v5

@@ 4 Homogeneous agents
    Suppose each (arm, public type) pair appears in at least $N_{\mathcal{P}}$ rounds of the warm-up stage, with probability $1$ over the agents’ beliefs. Here $N_{\mathcal{P}}<\infty$ is a parameter determined by the prior $\mathcal{P}$, as per Eq. (4.3). Then the mechanism is BIR over the main stage.
@@
$\mathtt{gap}(\mathcal{P}):=\textstyle\operatornamewithlimits{\mathbb{E}}_{\psi\sim\mathcal{P}}\left[\,U_{\psi}(a^{*}(\psi))-\tfrac{1}{n}\cdot\sum_{a\in[n]}U_{\psi}(a)\,\right],$ (4.2)
@@
$U_{\mathcal{P}}(\sigma^{\prime})$ $=\operatornamewithlimits{\mathbb{E}}_{\psi\sim\mathcal{P}}\left[\,(1-\epsilon+\epsilon/n)\cdot U_{\psi}(a^{*}(\psi))+\tfrac{\epsilon}{n}\cdot{\textstyle\sum_{a\neq a^{*}(\psi)}}U_{\psi}(a)\,\right]$
$\geq\operatornamewithlimits{\mathbb{E}}_{\psi\sim\mathcal{P}}\left[\,\sigma_{a^{*}(\psi)}(\psi)\cdot U_{\psi}(a^{*}(\psi))+\sigma_{a}(\psi)\cdot{\textstyle\sum_{a\neq a^{*}(\psi)}}U_{\psi}(a)\,\right]=U_{\mathcal{P}}(\sigma)\geq U_{\mathtt{out}}.$ (4.5)
The first inequality in (4.5) holds since $\sigma_{a}(\psi)\geq\epsilon/n$, so for each state the right-hand side shifts probabilities towards arms with weakly smaller expected utility. It follows that policy $\sigma^{\prime}$ is BIR. So, this policy gets within $\delta$ from the supremum in (4.4). This holds for any $\delta>0$. ∎
@@
$\mathbb{P}_{\mathtt{Bayes}}\left[\,U_{\psi}(\,a^{*}(\overline{\psi})\,)\geq U_{\psi}(\,a^{*}(\psi)\,)-2\delta\mid\psi^{*}=\psi\,\right]\geq 1-p_{\mathtt{err}}.$ (4.6)
@@
$U_{\mathcal{P}}\left(\,(\epsilon,\overline{\psi})\text{-$\mathtt{greedy}$}\,\right)$ $=\operatornamewithlimits{\mathbb{E}}\left[\,(1-\epsilon)\cdot U_{\psi}(a_{\overline{\psi}})+\tfrac{\epsilon}{n}\cdot{\textstyle\sum_{a\in[n]}}U_{\psi}(a)\,\right]$
$\geq\operatornamewithlimits{\mathbb{E}}\left[\,(1-\epsilon)\cdot U_{\psi}(a^{*}(\psi))+\tfrac{\epsilon}{n}\cdot{\textstyle\sum_{a\in[n]}}U_{\psi}(a)\,\right]-p_{\mathtt{err}}-2\delta$
$=\operatornamewithlimits{\mathbb{E}}\left[\,(1-2\epsilon)\cdot U_{\psi}(a^{*}(\psi))+\tfrac{2\epsilon}{n}\cdot{\textstyle\sum_{a\in[n]}}U_{\psi}(a)\,\right]+\epsilon\cdot\mathtt{gap}(\mathcal{P})-p_{\mathtt{err}}-2\delta$
$*\displaystyle\geq U_{\mathcal{P}}(\sigma^{*}),\quad\text{where we denote policy $(2\epsilon,\psi^{*})\text{-$\mathtt{greedy}$}$ with $\sigma^{*}$}.$
(All expectations are over $\psi\sim\mathcal{P}$.) The first inequality holds by (4.6) (and the fact that utilities are at most $1$). The last inequality holds because $\epsilon\cdot\mathtt{gap}(\mathcal{P})-p_{\mathtt{err}}-2\delta\geq 0$ by our choice of $\epsilon,\delta$.
@@ 6.2 Our results
$U^{*}_{\mathcal{P}}(\theta)$ $:=\operatornamewithlimits{\mathbb{E}}_{\psi\sim\mathcal{P}}\left[\,U_{\psi}\left(\,a^{*}(\psi,\theta),\,\theta\,\right)\,\right]$ (6.8)
@@
$U^{*}_{\mathcal{P}}(\theta,\theta_{\mathtt{rep}})$ $:=\operatornamewithlimits{\mathbb{E}}_{\psi\sim\mathcal{P}}\left[\,U_{\psi}\left(\,a^{*}(\psi,\theta^{\prime}),\,\theta\,\right)\,\right],\quad\text{where }\theta^{\prime}=(\theta_{\mathtt{pub}},\theta_{\mathtt{rep}}).$ (6.9)
@@ 8 Conclusions and open questions
Our work sets the stage for further investigation. First, we conjecture that “data-adaptive” mechanisms (i.e., mechanisms that can adapt to observations in an unrestricted way) cannot improve our guarantees on the estimation error, in the sense that all negative results (Theorems 4.5, 6.8 and 7.5) should essentially carry over to data-adaptive mechanisms. In particular, for homogeneous agents the key technical issue is to extend the lower bound on estimators (Lemma C.1) to allow for data-adaptivity. First, one could ask concrete questions that go slightly beyond our modelling approach. Can one match our guarantees for heterogeneous agents if the type space is not fully known, or the type frequencies are known very imprecisely? As the duration of the warm-up stage could be prohibitively large in practice, particularly in the presence of “difficult types”, can one reduce it and/or use some of its data for the counterfactual estimation? Second, one could ask a more “conceptual” question: what if the patients cannot fully understand or quantify their risk preferences, i.e., their own private types? What would be more realistic ways to elicit their preferences, and how to incorporate that into incentivized medical trials?
@@
[…]

### v6

@@ 4 Homogeneous agents
    Suppose each arm appears in at least $N_{\mathcal{P}}$ rounds of the warm-up stage, with probability $1$ over the agents’ beliefs. Here $N_{\mathcal{P}}<\infty$ is a parameter determined by the prior $\mathcal{P}$, as per Eq. (4.2). Then the mechanism is BIR over the main stage.
@@
$\mathtt{gap}(\mathcal{P}):=\textstyle\operatornamewithlimits{\mathbb{E}}_{\psi\sim\mathcal{P}}\left[\,U_{\psi}(a^{*}(\psi))-\tfrac{1}{n}\cdot\sum_{a\in[n]}U_{\psi}(a)\,\right]$
@@
The first inequality in (LABEL:eq:lm:pf:bench-single) holds since $\sigma_{a}(\psi)\geq\epsilon/n$, so for each state the right-hand side shifts probabilities towards arms with weakly smaller expected utility. It follows that policy $\sigma^{\prime}$ is BIR. So, this policy gets within $\delta$ from the supremum in (4.3). This holds for any $\delta>0$. ∎
@@
$\mathbb{P}_{\mathtt{Bayes}}\left[\,U_{\psi}(\,a^{*}(\overline{\psi})\,)\geq U_{\psi}(\,a^{*}(\psi)\,)-2\delta\mid\psi^{*}=\psi\,\right]\geq 1-p_{\mathtt{err}}$
@@
(All expectations are over $\psi\sim\mathcal{P}$.) The first inequality holds by (4) (and the fact that utilities are at most $1$). The last inequality holds because $\epsilon\cdot\mathtt{gap}(\mathcal{P})-p_{\mathtt{err}}-2\delta\geq 0$ by our choice of $\epsilon,\delta$.
@@ 6.2 Our results
$U^{*}_{\mathcal{P}}(\theta)$ $:=\operatornamewithlimits{\mathbb{E}}_{\psi\sim\mathcal{P}}\left[\,U_{\psi}\left(\,a^{*}(\psi,\theta),\,\theta\,\right)\,$
@@
$U^{*}_{\mathcal{P}}(\theta,\theta_{\mathtt{rep}})$ $:=\operatornamewithlimits{\mathbb{E}}_{\psi\sim\mathcal{P}}\left[\,U_{\psi}\left(\,a^{*}(\psi,\theta^{\prime}),\,\theta\,\right)\,\right],\quad\text{where }\theta^{\prime}=(\theta_{\mathtt{pub}},\theta_{\mathtt{rep}})$
@@ 8 Conclusions and open questions
Our work sets the stage for further investigation. First, one could ask concrete questions that go slightly beyond our modelling approach. Can one match our guarantees for heterogeneous agents if the type space is not fully known, or the type frequencies are known very imprecisely? As the duration of the warm-up stage could be prohibitively large in practice, particularly in the presence of “difficult types”, can one reduce it and/or use some of its data for the counterfactual estimation? Second, one could ask a more “conceptual” question: what if the patients cannot fully understand or quantify their risk preferences, i.e., their own private types? What would be more realistic ways to elicit their preferences, and how to incorporate that into incentivized medical trials?
@@
$\mathtt{LLR}_{[T_{0}]}\left(\,\psi^{*},\psi^{\prime}\,\right)>0\quad\text{for any other state $\psi^{\prime}\in\mathtt{support}(\mathcal{P})\setminus\left\{\,\psi^{*}\,\right\}$}.$ (A.1)